Oil flows from a tank through a 50 mm diameter pipe. If the average velocity is 1.3 m/s, determine how long it will take for 2500 kg of oil to flow out. Take S.G (oil) = 0.78


1
Expert's answer
2021-11-17T10:13:01-0500

given:

dia = 50mm

velocity = 1.3m/s

load = 2500kg

S.G = 780kg/m3

Volume of soil = load/S.G

= 2500/780

= 3.2 m3

Discharge = A x velocity

= /4(0.05)2 x 1.3

= 2.55 x 10-3m3/s

Time required = 3.2/2.55 x 10-3

   = 21 min